OpenStudy (anonymous):

45 is 30% of what number?

OpenStudy (anonymous):

=45/0.3 = 150 Answer: 150 Fraction's way: = 45/(30%/100%) = 45/(3/10) = 45(10/3) = 450/3 or 150

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20090922045843AAGg7aI this helps!

OpenStudy (anonymous):

yup it does...that is what i got this one from

OpenStudy (anonymous):

:)